25x to the second power plus 20x plus 4 equals 100
Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53427)

.
25x to the second power plus 20x plus 4 equals 100
~~~~~~~~~~~~~~~~~~~~~~~

25x^2 + 20x + 4 = 100.


Left side is a perfect square

    (5x+2)^2 = 100.


Hence, taking the square root

    5x + 2 = +/- 10.


Thus, one case is  5x + 2 = 10,  giving  5x = 10-2 = 8,     x = 8/5 = 1.6.

Another   case is  5x + 2 = -10, giving  5x = -10-2 = -12,  x = -12/5 = -2.4.

Solved completely.